OpenStudy (anonymous):

A truck driver slams on the brakes and the velocity of the truck changes from ‹11, 0, 0 › m/s to ‹ 8, 0, 0 › m/s in 6 seconds. The truck has a mass of 12000 kg. Compute the (average) net force exerted on the truck by the surroundings. What is the x component of the net force (in Newtons)?

OpenStudy (anonymous):

Quick Please !

OpenStudy (anonymous):

acceleration a=change in velocity/time =(8-11)/6=-0.5 m/s^2 so net force on the truck is F=ma=12000*(-0.5)=-6000 N x component of force is -6000N